Please show the detail, thank you!

(1 point) (a) Let -4 -7 -2 -4 V1 = and V2 = 1 6 0 2 and let W = span{V1, V2}. Apply the Gram-Schmidt procedure to vi and V2 t-13 2 (b) Consider the vector v = - Find V E W such that || V – v || is as small as possible. 15 8 V = I
Answer #1

(a) Vos -4 -2 1 V2= -4 6 2 K-1 K-1 W = span { ,,U2} By Gram-Schimidt procen, un= Vk - broj u;l = UK= Vk-Z <uj, Ok? uj so usFor any doubt please mention in comment box. Please like. Thanks